OpenStudy (anonymous):

1) y = sqrt(x), x = 1, Δx = 1 Δy= ? dy= ? 2) y = ex, x = 0, Δx = 0.5 Δy= ? dy= ?

OpenStudy (tkhunny):

If y = sqrt(x), then dy = dx/(2sqrt(x)) It's just the first derivative.

OpenStudy (caozeyuan):

when x=1,y=1,when x=x+deltax=2,y=sqrt(2), so delta y=sqrt(2)-1=0.414 approx

OpenStudy (caozeyuan):

for dy use the formula from tk

OpenStudy (caozeyuan):

and do the same for part 2